• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 18:30:40 on October 25, 2009

    Entry number 296658

    keyword=PD called, we request controlled access to reset level meters

    The PD called and relayed a message from Cryo that we are overloading the
    ESR with our magnet flows. It seems that we are pulling in too much
    cryogen to cool our magnets. This is similar to what we had yesterday: a
    stuck level meter while the level is at 100%. I discussed this with the
    tech-on-call an hour ago, but he suggested to wait until things stabilize
    (no significant change in the level has happened since), and has not been
    reachable. Since we want to avoid tripping the ESR (it is affecting
    other halls) I decide to go in and reset the level meters in both Left
    and Right arm, just as we did multiple times yesterday.

    I will not touch the flow meter on the Q1R A while I am down. It will
    therefore stay negative and prevent us from ramping up the Right arm
    until that is fixed.

    A copy of this log entry has been emailed to: paschke